A Carpernter used 26ft. of molding in three pieces to trim a garage door. If the long piece was 5 ft. longer than twice the length of each shorter piece, what is the length of the longest piece?
Answer by ptaylor(1894)

Let x=length of two shorter pieces
Then 2x+5=length of the long piece

Now we know that the length of all three pieces equals 26 ft so our equation to solve is:
x+x+(2x+5)=26
4x=21
x=5.25ft
2x+5=2(5.25)+5=10.50+5 =15.50ft for longest piece
Hope this helps----ptaylor
ck
15.50+10.50=26 ft
2x+5=2(5.25)+5=15.50ft
Hope this helps---ptaylor